# Towards Stabilized and Efficient Diffusion Transformers through Long-Skip-Connections with Spectral Constraints

### v4

@@ 4.4 Ablation studies
After incorporating LSCs into DiT, we can statically choose the timesteps for caching with feature stability, thus we can identify the dynamic phase by caching similarity shown in Figure 3. However, previous work has observed that the denoising process is split into different phases in U-Net diffusion models by Zhang et al. [49] and  Li et al. [13], highlighting the hidden dynamic phases which may lead to better performance. A heat map visualizing inner feature dynamics within blocks is shown in Figure 5, and significant changes are observed in the early denoising phase. In Table 5, we further validate this observation: under equivalent throughput, caching in the later timesteps (700$\rightarrow$50), outperforms caching in the earlier timesteps (950$\rightarrow$300), achieving superior PSNR, LPIPS, and SSIM scores. Additionally, we segmented the rapidly changing timesteps and experimented with additional caching ranges. All of these results show that increasing the ratio of the former phase significantly improves caching performance.
Table 5: Ablation study on different timestep ranges for caching. Total timesteps is 1000. We only perform cache in Skip-DiT  at caching timesteps. We highlight the best metrics in bold.
@@
Table 6: Compatibility of Skip-DiT  with other caching methods. The performance or other caching methods on vanilla DiT are shown in the table 2. Obvious improvements can be seen especially in PAB. We highlight the best metrics in bold.
Method $\Delta$VBench (%) $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$
@@
LSC Cached Similarity(%$\uparrow$ VBench(%) $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$
@@
NFE Model $\Delta$VBench(%) PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$
@@
5 Conclusion
In this work, we identify the Dynamic Feature Instability in vanilla DiT models and introduce Skip-DiT , a Long-Skip-Connections enhanced DiT model to produce more stable features and build training and inference efficient diffusion transformers. We designed a simpler but more efficient caching mechanism for Skip-DiT  to improve inference efficiency in video and image generation tasks. By improving the stability of dynamic features, Skip-DiT  unlocks the potential to cache most blocks while maintaining high generation quality. In addition, we guarantee stability and efficiency with theoretical analysis from the perspective of the spectral norm. Our proposed DiT achieves significant speedup in training convergence and model capability. With the caching mechanism, we achieve maximum speedup in cache-based acceleration while maintaining the highest similarity to the original outputs. Furthermore, we identify the dynamic phase of Skip-DiT  for a better caching strategy, demonstrating the effectiveness of caching. We also show that Skip-DiT  is compatible with other caching methods, further extending its applicability. Overall, Skip-DiT  can be seamlessly integrated with various DiT backbones, enabling real-time, high-quality video and image generation, while consistently outperforming baseline methods. We believe that Skip-DiT  provides a simple yet powerful foundation for advancing future research and applications.
6 Limitations
Acknowledgement
Supplementary Material
@@ 7.1 Theorem1
Define layer transformations for $L/2<l\leq L$ and their Jacobian matrices, where $\mathcal{T}$ denotes the Transformer block and $10<\alpha<$ 1:
@@ 8 Class-to-image Generation Experiments
Peebles and Xie [22] proposed the first diffusion model based on the transformer architecture, and it outperforms all prior diffusion models on the class conditional ImageNet [7] 512$\times$512 and 256$\times$256 benchmarks. We add skip connections to its largest model, DiT-XL, to get Skip-DiT. We train Skip-DiT on class conditional ImageNet with resolution 256$\times$256 from scratch with completely the same experimental settings as DiT-XL, and far exceeds DiT-XL with only around 38% of its training cost.
@@
Methods FID $\downarrow$ sFID $\downarrow$ IS $\uparrow$ PrecisionRecallSpeedup
@@
DiT-XL 2.30 4.71 276.26 82.68 57.65 1.00$\times$
FORA 2.45 5.44 265.94 81.21 58.36 1.57$\times$
Delta-DiT 2.47 5.61 265.33 81.05 58.83 1.45$\times$
Faster-Diff 4.96 10.19 223.21 75.21 59.28 1.42$\times$
@@
DiT-XL 9.49 7.17 122.49 66.66 67.69 1.00$\times$
FORA 11.72 9.27 113.01 64.46 67.69 1.53$\times$
Delta-DiT 12.03 9.68 111.86 64.57 67.53 1.42$\times$
Faster-Diff 22.98 18.09 80.41 56.53 67.41 1.42$\times$
@@
Skip-DiT 8.37 6.50 127.63 68.06 67.89 1.00$\times$
$n=2$ 9.25 7.09 123.57 67.32 67.40 1.46$\times$
$n=3$ 10.18 7.72 119.60 66.53 67.84 1.71$\times$
$n=4$ 11.37 8.49 116.01 65.73 67.32 1.92$\times$
We evaluate Skip-DiT  and compare its performance against two other caching methods: $\Delta$-DiT and FORA. As shown in Table 10, Skip-DiT achieves a 1.46$\times$ speedup with only a minimal FID loss of $0.02$ when the classifier-free guidance scale is set to 1.5, compared to the 7–8$\times$ larger losses observed with $\Delta$-DiT and FORA. Moreover, even with a 1.9$\times$ acceleration, Skip-DiT  performs better than the other caching methods. These findings further confirm the effectiveness of Skip-DiT  for class-to-image tasks.

### v5

@@ 4.4 Ablation Studies
[…]
Table 5: Ablation study on different timestep ranges for caching. Total timesteps is 1000. Caching is performed at caching timesteps.
@@
Table 6: Compatibility of Skip-DiT with other caching methods. Obvious improvements can be seen especially in PAB.
[…]
@@
5 Conclusion
In this work, we identify the Dynamic Feature Instability in vanilla DiT models and introduce Skip-DiT , a Long-Skip-Connections enhanced DiT model to produce more stable features and build training and inference-efficient diffusion transformers. We designed a simple, static, but more efficient caching mechanism for Skip-DiT to improve inference efficiency in video and image generation tasks. By improving the stability of dynamic features, Skip-DiT unlocks the potential to cache most blocks while maintaining high generation quality. In addition, we guarantee stability and efficiency with theoretical analysis from the perspective of the spectral norm and visualizations. Furthermore, we identify the dynamic phase of Skip-DiT for a better caching strategy, demonstrating the effectiveness of caching. We also show that Skip-DiT is compatible with other caching methods, further extending its applicability. Overall, we believe that Skip-DiT provides a simple yet powerful foundation for advancing future research and applications.
Acknowledgement
Supplementary Material
6 Limitations
7 Supplementary Experiments
@@
[…]
@@
[…]
@@ 8 Detailed Proof of Theorems
  • Transformer Blocks Homogeneity: Identical noise reduction ratio across layers $\gamma_{l}=\gamma,\qquad\forall l\in\{1,...,L\}$. Thus, the complete model satisfies
@@ 8.1 Theorem1
Define layer transformations for $L/2<l\leq L$ and their Jacobian matrices, where $\mathcal{T}$ denotes the Transformer block and $0<\alpha<1$:
@@ 9 Class-to-image Generation Experiments
Peebles and Xie [23] proposed the first diffusion model based on the transformer architecture, and it outperforms all prior diffusion models on the class conditional ImageNet [7] $\times512$ and $\times256$ benchmarks. We add skip connections to its largest model, DiT-XL, to get Skip-DiT. We train Skip-DiT on class conditional ImageNet with resolution $\times256$ from scratch with completely the same experimental settings as DiT-XL, and far exceeds DiT-XL with only around 38% of its training cost.
@@
Methods FID $\downarrow$ sFID $\downarrow$ IS $\uparrow$ Precision% Recall% Speedup
@@
[…]
@@
[…]
@@
[…]